OpenStudy (anonymous):

What is the simplified sum of √3+√12=

OpenStudy (anonymous):

3 square root 3

OpenStudy (campbell_st):

well look at the 2nd term and apply some laws for radicals \[\sqrt{12} = \sqrt{4 \times 3} = \sqrt{4} \times \sqrt{3}\] so the question can be written as \[\sqrt{3} + 2\sqrt{3} = \] hope it helps

OpenStudy (mandre):

√3+2√3= 3√3

OpenStudy (anonymous):

ya